Given a random variable with binomial distribution use moment generating technique to determine the mean


1
Expert's answer
2022-04-17T10:18:08-0400

The moment generating function for binomial distribution:

"f_x(t)=\Epsilon_x(e^{tx})=\sum_{k=0}^{n}e^{tk}C_n^kp^k(1-p)^{n-k}=\\
=\sum_{k=0}^{n}C_n^k(pe^t)^k(1-p)^{n-k}=(pe^t+(1-p))^n\\
f_x^{(1)}(t)=n(pe^t+(1-p))^{n-1}pe^t"

Mean is the first central moment and is calculated by the next formula:

"\mu=f_x^{(1)}(0)=n(pe^0+(1-p))^{n-1}pe^0=np"